# Three-dimensional fracton topological orders with boundary Toeplitz braiding

### v2

@@ V Conclusions
In this paper, we constructed a type of 3D topological fracton topological orders whose surface states support a nontrivial form of braiding statistics, termed Toeplitz braiding. This boundary phenomenon arises due to the block-tridiagonal Toeplitz structure of the $K$-matrix in the infinite-component Chern-Simons theory. This phenomenon represents a significant departure from conventional braiding statistics observed in 2D topological orders. Here, we further summarize the paper in details. First, in the field of free-fermion topological insulators, boundary zero modes of Toeplitz matrices play a vital role. In this paper, we substantially extend the topological physics of boundary zero modes of Toeplitz matrices to the present strongly correlated systems. Second, our research categorizes nontrivial $K$-matrices into two distinct types based on their boundary zero modes. This classification provides a deeper understanding of the underlying mathematical structures and their physical implications in 3D fracton topological orders. Third, we conduct extensive numerical simulations and finite-size scaling analyses to confirm the theoretical predictions. These simulations demonstrate the robustness of Toeplitz braiding.
@@
There are several future directions. First of all, it will be interesting to construct the lattice model realization of iCS field theories that supports boundary Toeplitz braiding, which may pave the way for further numerical or experimental study on Toeplitz braiding. From the perspective of fractional quantum Hall states or spin liquid, the stacking process described in Sec. II.1 can be realized by condensing topological excitations followed by adding interlayer interaction of condensation currents or spin-spin interaction [43, 44, 15, 12], which may be also useful here. Detailed construction that explicitly enables Toeplitz braiding will be left to future discussion. Furthermore, following the construction of topologically-ordered non-liquid states in this paper, the construction of higher-dimensional topologically-ordered non-liquid states is directly allowed. For instance, by stacking (3+1)D topological orders [45, 46, 47, 48], we can obtain a (4+1)D topologically-ordered non-liquid state. Similarly, stacking (2+1)D topological orders in two independent spatial dimensions can also yield (4+1)D topologically-ordered non-liquid states. The study of these theories and the duality between them is left for future research. We can also compute entanglement entropy via partitioning the system into two subsystems along $z$ directions and study whether and how the presence of Toeplitz braiding leaves its fingerprint in the scaling behavior of entanglement entropy.

### v4

@@ V Conclusions
In this paper, we constructed a type of 3D topological fracton topological orders whose surface states support a nontrivial form of braiding statistics, termed Toeplitz braiding. This boundary phenomenon arises due to the block-tridiagonal Toeplitz structure of the $K$-matrix in the infinite-component Chern-Simons theory. This phenomenon represents a significant departure from conventional braiding statistics observed in 2D topological orders. Here, we further summarize the paper in details. First, in the field of free-fermion topological insulators, boundary zero modes of Toeplitz matrices play a vital role. In this paper, we substantially extend the topological physics of boundary zero modes of Toeplitz matrices to the present strongly correlated systems. Second, our research categorizes nontrivial $K$-matrices into two distinct types based on their boundary zero modes. This classification provides a deeper understanding of the underlying mathematical structures and their physical implications in 3D fracton topological orders. Third, we conduct extensive numerical simulations and finite-size scaling analyses to confirm the theoretical predictions. These simulations demonstrate the robustness of Toeplitz braiding. Fourth, we also provide a discussion on symmetry fractionalization in iCS field theories, concluding that the boundary zero modes of $K$ influence the fractionalization pattern.
@@
There are several future directions. First of all, the categorization of iCS field theories with Toeplitz braiding is primarily for mathematical convenience. How to physically distinguish Type-I and Type-II iCS field theories is left for future research. It will also be interesting to construct the lattice model realization of iCS field theories that supports boundary Toeplitz braiding, which may pave the way for further numerical or experimental study on Toeplitz braiding. From the perspective of fractional quantum Hall states or spin liquid, the stacking process described in Sec. II.1 can be realized by condensing topological excitations followed by adding interlayer interaction of condensation currents or spin-spin interaction [44, 45, 15, 12], which may be also useful here. Detailed construction that explicitly enables Toeplitz braiding will be left to future discussion. We might also consider enlarging blocks $A$ and $B$ in the Toeplitz $K$-matrix, which might affect the number of boundary zero modes and reshape the braiding statistics. Breaking the strict Toeplitz character of the $K$-matrix and considering the “stability” of braiding statistics against stacking faults is also an intriguing topic. Furthermore, following the construction of topologically-ordered non-liquid states in this paper, the construction of higher-dimensional topologically-ordered non-liquid states is directly allowed. For instance, by stacking (3+1)D topological orders [46, 47, 48, 49], we can obtain a (4+1)D topologically-ordered non-liquid state. Similarly, stacking (2+1)D topological orders in two independent spatial dimensions can also yield (4+1)D topologically-ordered non-liquid states. The study of these theories and the duality between them is left for future research. We can also compute entanglement entropy via partitioning the system into two subsystems along $z$-direction and study whether and how the presence of Toeplitz braiding leaves its fingerprint in the scaling behavior of entanglement entropy.